OpenStudy (anonymous):

what is the sin of A A.8/15 B.8/17 C.17/15 D.15/17

OpenStudy (anonymous):



OpenStudy (anonymous):

Because \(\Large AC^2=17^2=289\) and \(\Large AB^2 + BC^2=8^2+15^2=289\) So, \(\Large \triangle ABC \) is a right triangle. \(\Large \Rightarrow sinA=\frac{BC}{AC}=\frac{15}{17}\)

OpenStudy (anonymous):

thank you will you help me some more